The pOH of a solution is found to be 9. What is the hydrogen ion ( H1+) concentration for this solution?

a) 1x10^-5

b) 1x10^-9

c) 1x10^-3

d) 1x10^-12
1
Expert's answer
2014-06-18T08:40:40-0400
pH + pOH = 14
pH = 14 - pOH = 5
pH = -lg [H+]
[H+] = 10-pH = 1x10-5
So, the correct option is (a): 1x10-5